OpenStudy (anonymous):

What is the simplified form of (15x^8/24y^5) divided by (4x^4/8x^2)

OpenStudy (klimenkov):

\[\frac {15 x^8}{24y^5}\div\frac{4x^4}{8x^2}=\frac {15 x^8}{24y^5}\cdot\frac{2}{x^2}=\frac {15 x^6}{12y^5}=\frac{15}{12}x^6y^{-5}\]